# Terrain Diffusion: A Diffusion-Based Successor to Perlin Noise in Infinite, Real-Time Terrain Generation

### v2

@@ 2. Related Works
Learned terrain generation. GAN-based terrain models (Goodfellow et al., 2020; Voulgaris et al., 2021; Spick and Walker, 2019; Beckham and Pal, 2017; Argudo et al., 2018) can generate convincing local relief, but they operate on fixed crops and do not tile, limiting them to bounded worlds. Diffusion-based synthesis (Hu et al., 2024; Guérin et al., 2017; Borne-Pons et al., 2025) further improves fidelity and control, but also assumes finite canvases and requires relatively significant compute. Jain et al. (Jain et al., 2023) is closest to our work, offering infinite terrain generation by sampling diffusion-based tiles and blending them with a Perlin-based kernel. Because tiles are generated independently and the kernel has no awareness of broader context, structure remains tied to perlin noise rather than the learned model. In contrast, Terrain Diffusion couples all tiles through a shared global context and fuses tiles through a fully learned, context-aware mechanism. Procedural noise is used only for defining continental layouts, where data is sparse and simple enough that more complex alternatives would provide little benefit while reducing user control.
MultiDiffusion and unbounded generation. Several works extend generative models beyond image bounds seen in training. InfinityGAN (Lin et al., 2022) produces infinite images with GANs but does not extend to diffusion models, limiting scalability. MultiDiffusion (Bar-Tal et al., 2023) and Mixture of Diffusers (Jiménez, 2023) generate images larger than the training canvas but still assume a bounded final extent. BlockFusion (Wu et al., 2024) and WorldGrow (Li et al., 2025) generate worlds by conditioning each tile on its neighbors, producing continuous worlds but without seed consistency, since outputs depend on sampling order. In contrast, Terrain Diffusion defines a seed-consistent InfiniteDiffusion algorithm whose outputs are order invariant and allow constant-time random-access generation over an infinite domain.
@@ 3.1. A Review of MultiDiffusion
To accomplish this, MultiDiffusion defines $n$ windows indexed by $i\in[n]$. In the finite setting, a region $R$ is any rectangular subset of the $H^{\prime}\times W^{\prime}$ coordinate grid, while a window region has a fixed size $H\times W$. In MultiDiffusion, each window $i$ is assigned a window region $R_{i}$. For an image $J\in\mathcal{J}$, we write $J[R]\in\mathcal{I}$ for the values of $J$ on the coordinates in $R$.
@@ 3.3. Practical Querying of InfiniteDiffusion
To make queries practical, we avoid recomputing the same window updates across recursive calls. Instead, for each image $J_{t}$ we maintain two corresponding infinite tensors: $A_{t}$, which stores the numerator in Eq.  2, and $B_{t}$, which stores the denominator. Both tensors are initialized to zero. These infinite tensors are typically stored in tiles, which can be loaded and unloaded as necessary to conserve memory. When a query $J_{t}[R]$ occurs, we identify all the windows required to generate the region, and process all previously unprocessed windows, populating the desired regions of $A_{t}$ and $B_{t}$. Then $J_{t}[R]=A_{t}[R]/B_{t}[R]$. Final generation proceeds as a recursive process that begins by sampling $J_{T}$ as Gaussian noise. $J_{0}[R]$ is obtained by recursively applying the query routine at all earlier steps. In summary, the routine below computes $J_{t-1}[R]$ by evaluating only the windows that overlap $R$, caching each window’s contribution in $A_{t-1}$ and $B_{t-1}$ so that future queries can reuse the same results. The tensors $A_{t-1}$, $B_{t-1}$, and the set of processed windows $P$ are mutated in-place.
Algorithm 1 Querying $J_{t-1}[R]$ with InfiniteDiffusion. See Fig.  2
@@
  $A_{t-1}[R_{i}]\leftarrow A_{t-1}[R_{i}]+W_{i}\otimes\Phi(J_{t}[R_{i}]\mid y_{i})$
  $B_{t-1}[R_{i}]\leftarrow B_{t-1}[R_{i}]+W_{i}$
@@
$P\leftarrow P\cup\kappa(R)$
Output: $J_{t-1}[R]=A_{t-1}[R]/B_{t-1}[R]$
@@ 3.4. Tractability via Truncated $T$
There is one final but critical barrier in making InfiniteDiffusion practical. Each query $J_{t-1}[R]$ typically requires a region of $J_{t}$ larger than $R$ itself, as pixels near the edge of $R$ are generated by windows that extend beyond $R$. In the sliding-window case, regions grow quadratically in area, yielding an overall $O(T^{3})$ time complexity for final generation. For traditional MultiDiffusion, which relies on continuously fusing diffusion paths over dozens of steps, this complexity renders infinite generation effectively impossible.
@@
Crucially, we find that our framework retains striking coherence even when generation is aggressively truncated to $T=1$ or $2$. We hypothesize that because global structure is largely determined from hierarchical conditioning, overlapping windows only need to enforce local consistency that is already largely aligned through shared noise. Additionally, because the models are trained on random crops, they learn an approximately translation invariant representation that bases predictions on local patterns, which are shared across overlapping windows, rather than absolute position, further reducing seams. Together, these factors make the method robust even under minimal iterative refinement.
@@ 3.5. Properties of InfiniteDiffusion
Formal proofs for all properties are provided in appendix   B.
@@
A central property of InfiniteDiffusion is that it preserves the seed-consistent behavior that makes procedural noise functions useful for procedural world generation. Seed consistency follows directly from how the process is defined. Once a seed is fixed, the initial noise image $J_{T}$ is completely determined. The query routine is a deterministic function of its inputs, so $J_{t-1}$ is a deterministic function of $J_{t}$, and caching only memoizes intermediate results of this deterministic computation. This argument requires that the conditioning variables $y_{i}$ are themselves seed-consistent, which is typically trivial when $y_{i}$ is constant or generated by the same InfiniteDiffusion algorithm, as in this work. By composing these steps, every image $J_{t}$ and every region $J_{t}[R]$ becomes a fixed function of the original seed. In particular, requesting the same region again, or in a different order relative to other regions, always returns exactly the same value, since the computation depends only on the seed and the region, not on the sequence of queries.
@@ 3.6. An Open Source Infinite Tensor Framework for Unbounded Inference
To support unbounded generation without exceeding memory limits, we introduce the Infinite Tensor framework, a Python library that enables sliding window computation over tensors with infinite dimensions. It allows models to process arbitrarily large images as if they were standard PyTorch tensors while keeping only the visible region in memory. Each operation is performed through a fixed-sized sliding window that dynamically loads and evicts data as sampling progresses, permitting inference on arbitrarily large scenes with bounded memory use.
@@ 4. Data
To enable truly global generation, we construct a seamless global elevation dataset merging land topography from the 90m MERIT DEM (Yamazaki et al., 2017) and ocean bathymetry from ETOPO1 (NOAA National Geophysical Data Center, 2009), supplemented by climatic data from WorldClim (Fick and Hijmans, 2017). We process this data into equal-area tiles by dynamically stretching longitude based on latitude; this ensures that pixel sizes represent a consistent physical area, allowing the diffusion model to learn features with minimal polar distortion. The dataset is split into $\times2048$ training tiles, with specific details on coastline smoothing, bathymetric merging, and sampling heuristics provided in Appendix  D.
@@ 5.1. Signed Square-Root Transform
Terrain tiles vary significantly in elevation range. Under a fixed diffusion noise schedule, this leads to uneven effective SNR: low-relief regions behave as though exposed to stronger noise, while high-relief terrain is affected much less. In the raw elevation space, tiles with higher absolute elevations exhibit larger variance. To reduce this variation, we apply a signed square-root transform $z\mapsto\mathrm{sign}(z)\sqrt{|z|}$. The transform compresses high-relief values and distributes variance across tiles more uniformly. Namely, the correlation between the mean and log standard deviation of tiles decreases from $0.66$ to $0.31$. In practice, this allows us to train with a more focused noise distribution and enhances the visibility of small features, especially coastlines. Additional details in Appendix  E.
@@ 5.2. Stabilization Via Laplacian Encodings
Due to normalization, the large dynamic range of Earth elevations make model errors deceptively large in absolute units. Even relatively small errors of $\sigma=0.01$ can correspond to $\pm 25$m noise after denormalization. While this noise is less obvious in images, it can become very apparent in interactive settings, where the noise is potentially much larger than the viewer or the surrounding scenery. To mitigate this, we predict a Laplacian-based representation comprising a low-frequency component, obtained by downsampling and blurring the original image, and a residual/high-frequency component given by subtracting the upsampled low frequency component from the original image.
@@ 5.3. A Hierarchical Model
Planetary terrain spans several orders of magnitude in scale, from continental structure to meter-level detail, making one-pass generation infeasible. Several previous works (Sharma et al., 2024; Zhang et al., 2025; Lee et al., 2023; Du et al., 2024; Zhou and Tang, 2024) have shown that MultiDiffusion produces incoherent and repetitive results when poorly conditioned, and InfiniteDiffusion does not natively provide a solution to this. We therefore organize generation into a small hierarchy of models operating at progressively finer resolutions. Each stage refines and conditions on the one above, maintaining large-scale coherence while producing realistic local detail. All models share a common EDM2 (Karras et al., 2024b) backbone with the modifications proposed in sCM (Lu and Song, 2025). The hierarchy begins with a coarse planetary model, which generates the basic structure of the world from a rough, procedural or user-provided layout. The next stage is the core latent diffusion model (Rombach et al., 2021), which transforms that structure into realistic 46km tiles in latent space. Finally, a consistency decoder expands these latents into a high-fidelity elevation map. We visualize the coarse-to-fine pipeline in Figure  3.
@@
The core latent diffusion model synthesizes 512×512 patches at a 90m resolution in signed-sqrt space, corresponding to 46km tiles. It predicts a 64x64 low-frequency elevation channel and latent map that compactly represents the corresponding residual. To supply the latent codes, we train a separate VAE-style autoencoder that shares the same U-Net backbone but omits diffusion-specific components. The model is optimized using L1 and LPIPS losses with a weak KL term to prevent overfitting. After training, the encoder processes each 2048×2048 tile in the dataset as a whole, and the resulting latent image is precomputed and stored alongside the tile. To maximize local quality, the imprecise VAE decoder is discarded, and a diffusion decoder learns to expand these latents into realistic and high-resolution residuals. During training of all models, we draw random crops from the latent image to learn a nearly translation-invariant representation, reflecting the fact that generation should be independent of absolute location. Conditioning is implemented by concatenating nearest-neighbor interpolated latents to the noisy input image at each diffusion step. We train both the autoencoder and diffusion decoder on 128×128 crops, but find the models generalize well when applied to higher-resolution patches.
@@
To facilitate long-range global coherence, the core model is conditioned on 4x4 patches of elevation data. Each pixel of the patch is about 23km, with the model prediction corresponding to the 2x2 interior. Each patch contains 3 channels: the mean elevation of the pixel, the 5th percentile elevation of the pixel, and a binary mask indicating which pixels have data available. We also provide the model with the tile’s mean temperature, temperature variation, annual precipitation, and precipitation seasonality for additional coherence and user control. Since climatic data is not available in the ocean, we replace missing values with a standard gaussian, ensuring the model accepts any combination of climatic variables in ocean regions.
@@ 5.4. Real-Time Planetary Scale Synthesis
While the 46 km tiles are realistic in isolation, large-scale coherence requires conditioning on broader planetary context. To provide this, we introduce a compact coarse diffusion model that generates the channels required for conditioning the core diffusion model. The user provides initial maps for these variables using hand-drawn sketches or procedural noise, which we found works as well as learned methods. During inference, these inputs are corrupted with gaussian noise according to the user’s preference on a per-channel basis, and concatenated against the usual diffusion inputs. The model follows the EDM2 design but with no downsampling or upsampling operations. This limits the receptive field by design, preventing the model from drifting toward the massive continental structures present in Earth data and avoiding conflicts where global priors override user guidance while still allowing strong local corrections.
@@ 6. Results
We evaluate Terrain Diffusion on visual fidelity and latency. All experiments use a single NVIDIA RTX 3090 Ti GPU. InfiniteDiffusion uses $T=2$ and a stride of 32 in the core model, and $T=1$ with a batch size of 1 in the coarse and decoder models. We use a separable linear weight window that is 1 at the window center and decreases linearly to a small value $\epsilon$ at the boundary, which reduces FID from $19.32$ with a constant map to $14.78$. The decoder model uses windows of size 512 with strides 384. Batch size is 16 for the core model and 1 for other components.
@@ 7. Discussion
The main limitation of our method is that each model relies on conditioning from the level above it, and global coherence deteriorates if this sequence is broken. Some top level must therefore be specified externally. In this work, simple Perlin noise is sufficient because continental-scale structure is extremely coarse, allowing Perlin to provide a reasonable starting point while remaining seed-consistent and controllable. In domains where a procedural prior is unavailable, a learned generator such as InfinityGAN (Lin et al., 2022) may provide a viable solution. Since our hierarchy can refine and upsample a coarse signal by several orders of magnitude, the traditional limitations of GANs are largely mitigated, as they only need to provide the coarsest global gradients. We experimented with an end-to-end hierarchy using a small padding-free GAN as the base model, which preserved the functional properties of our system, but it offered reduced controllability compared to noise-based alternatives. This direction remains open for future work.

### v3

@@ 2. Related Works
Learned terrain generation. GAN-based terrain models (Goodfellow et al., 2020; Voulgaris et al., 2021; Spick and Walker, 2019; Beckham and Pal, 2017; Argudo et al., 2018) can generate convincing local relief, but they operate on fixed crops and do not tile, limiting them to bounded worlds. Diffusion-based synthesis (Hu et al., 2024; Guérin et al., 2017; Borne-Pons et al., 2025) further improves fidelity and control, but also assumes finite canvases and requires relatively significant compute. Jain et al. (Jain et al., 2023) offers infinite terrain generation by sampling diffusion-based tiles and blending them with a Perlin-based kernel. Because tiles are generated independently and the kernel has no awareness of broader context, structure remains tied to perlin noise rather than the learned model. In contrast, Terrain Diffusion couples all tiles through a shared global context and fuses tiles through a fully learned, context-aware mechanism. Procedural noise is used only for defining continental layouts, where data is sparse and simple enough that more complex alternatives would provide little benefit while reducing user control.
MultiDiffusion and unbounded generation. Several works extend generative models beyond image bounds seen in training. InfinityGAN (Lin et al., 2022) produces infinite images with GANs but does not extend to diffusion models, limiting scalability. MultiDiffusion (Bar-Tal et al., 2023) and Mixture of Diffusers (Jiménez, 2023) generate images larger than the training canvas but still assume a bounded final extent. Auto-regressive methods such as BlockFusion (Wu et al., 2024) and WorldGrow (Li et al., 2025) generate worlds by conditioning each tile on its neighbors, producing continuous worlds but without seed consistency, since outputs depend on sampling order. Additionally, tiles must be produced sequentially, preventing efficient random access. In contrast, we define a seed-consistent InfiniteDiffusion algorithm whose outputs are order invariant and allow constant-time random-access generation over an infinite domain.
@@ 3.1. A Review of MultiDiffusion
To accomplish this, MultiDiffusion defines $n$ windows indexed by $i\in[n]$. In the finite setting, a region $R$ is any rectangular subset of the $H^{\prime}\times W^{\prime}$ coordinate grid, while a window region has a fixed size $H\times W$. In MultiDiffusion, each window $i$ is assigned a window region $R_{i}$. For an image $J\in\mathcal{J}$, we write $J[R]$ for the values of $J$ on the coordinates in $R$.
@@ 3.3. Practical Querying of InfiniteDiffusion
To make queries practical, we avoid recomputing the same window updates across recursive calls. Instead, for each image $J_{t}$ we maintain two corresponding infinite tensors: $A_{t}$, which stores the numerator in Eq.  2, and $B_{t}$, which stores the denominator. Both tensors are initialized to zero. These infinite tensors are typically stored in tiles, which can be loaded and unloaded as necessary to conserve memory. When a query $J_{t}[R]$ occurs, we identify all the windows required to generate the region, and process all previously unprocessed windows, populating the desired regions of $A_{t}$ and $B_{t}$. Then $J_{t}[R]=A_{t}[R]/B_{t}[R]$. Final generation proceeds as a recursive process that begins by sampling $J_{T}$ as Gaussian noise. $J_{0}[R]$ is obtained by recursively applying the query routine at all earlier steps. In summary, Algorithm  1 below computes $J_{t}[R]$ by evaluating only the windows that overlap $R$, caching each window’s contribution in $A_{t}$ and $B_{t}$ so that future queries can reuse the same results. The tensors $A_{t}$, $B_{t}$, and the set of processed windows $P_{t}$ are mutated in-place. See Figure  2 for a visualization.
Algorithm 1 Querying $J_{t}[R]$ ($t$ ¡ $T$) with InfiniteDiffusion.
@@
  $A_{t}[R_{i}]\leftarrow A_{t}[R_{i}]+W_{i}\otimes\Phi(J_{t+1}[R_{i}]\mid y_{i})$
  $B_{t}[R_{i}]\leftarrow B_{t}[R_{i}]+W_{i}$
@@
$P_{t}\leftarrow P_{t}\cup\kappa(R)$
Output: $J_{t}[R]=A_{t}[R]/B_{t}[R]$
@@ 3.4. Tractability via Truncated $T$
There is one final but critical barrier in making InfiniteDiffusion practical. Each query $J_{t}[R]$ typically requires a region of $J_{t+1}$ larger than $R$ itself, as pixels near the edge of $R$ are generated by windows that extend beyond $R$. In the sliding-window case, regions grow quadratically in area, yielding an overall $O(T^{3})$ time complexity for final generation. For traditional MultiDiffusion, which relies on continuously fusing diffusion paths over dozens of steps, this complexity renders infinite generation effectively impossible.
@@
Crucially, we find that our framework retains striking coherence even when generation is aggressively truncated to $T=1$ or $2$. We hypothesize that because global structure is largely determined from hierarchical conditioning, overlapping windows only need to enforce local consistency that is already largely aligned through shared noise. Additionally, our models are trained on random crops, so they learn an approximately translation invariant representation that bases predictions on local patterns, which are shared across overlapping windows, rather than absolute position, further reducing seams. Together, these factors make the method robust even under minimal iterative refinement.
@@ 3.5. Properties of InfiniteDiffusion
Formal proofs for all properties are provided in Appendix  B
@@
A central property of InfiniteDiffusion is that it preserves the seed-consistent behavior that makes procedural noise functions useful for procedural world generation. Seed consistency follows directly from how the process is defined. Once a seed is fixed, the initial noise image $J_{T}$ is completely determined. The query routine is a deterministic function of its inputs, so $J_{t}$ is a deterministic function of $J_{t+1}$, and caching only memoizes intermediate results of this deterministic computation. This argument requires that the conditioning variables $y_{i}$ are themselves seed-consistent, which is typically trivial when $y_{i}$ is constant or generated by the same InfiniteDiffusion algorithm, as in this work. By composing these steps, every image $J_{t}$ and every region $J_{t}[R]$ becomes a fixed function of the original seed. In particular, requesting the same region again, or in a different order relative to other regions, always returns exactly the same value, since the computation depends only on the seed and the region, not on the sequence of queries.
@@ 3.6. An Open Source Infinite Tensor Framework for Unbounded Inference
To support unbounded generation without exceeding memory limits, we introduce the Infinite Tensor framework [^1] Available at https://github.com/xandergos/infinite-tensor, a Python library that enables sliding window computation over tensors with infinite dimensions. It allows querying of arbitrarily large images as if they were standard PyTorch tensors while keeping only the visible region in memory. Each operation is performed through a fixed-sized sliding window that dynamically loads, evicts, or deletes data as sampling progresses. While Algorithm  1 theoretically requires unbounded storage, our framework can optionally cache operations $\mathcal{T}[R_{i}]\leftarrow\mathcal{T}[R_{i}]+x$ directly, so $\mathcal{T}$ is stored implicitly as a series of summations. Consequently, each contribution $x$ for any window need not be persistently stored; it can be deleted and simply recomputed on demand, permitting bounded disk use.
@@ 4. Data
To enable truly global generation, we construct a seamless global elevation dataset merging land topography from the 90m MERIT DEM (Yamazaki et al., 2017) and ocean bathymetry from ETOPO1 (NOAA National Geophysical Data Center, 2009), supplemented by climatic data from WorldClim (Fick and Hijmans, 2017). We process this data into equal-area tiles by dynamically stretching longitude based on latitude; this ensures that pixel sizes represent a consistent physical area, allowing the diffusion model to learn features with minimal polar distortion. The dataset is split into $\times2048$ training tiles, with specific details on coastline smoothing, bathymetric merging, and sampling heuristics provided in Appendix  E.
@@ 5.1. Signed Square-Root Transform
Terrain tiles vary significantly in elevation range. Under a fixed noise distribution, this leads to uneven effective SNR: low-relief regions behave as though exposed to stronger noise, while high-relief terrain is affected much less. In the raw elevation space, tiles with higher absolute elevations exhibit larger variance. To reduce this variation, we apply a signed square-root transform $z\mapsto\mathrm{sign}(z)\sqrt{|z|}$. The transform compresses high-relief values and distributes variance across tiles more uniformly. Namely, the correlation between the mean and log standard deviation of tiles decreases from $0.66$ to $0.31$. In practice, this allows us to train with a more focused noise distribution and enhances the visibility of small features, especially coastlines. Additional details in Appendix  F.
@@ 5.2. Stabilization Via Laplacian Encodings
Due to normalization, the large dynamic range of Earth elevations make model errors deceptively large in absolute units. Even relatively small errors of $\sigma=0.01$ can correspond to $\pm 25$m noise after denormalization. To mitigate this, we predict a Laplacian-based representation comprising a low-frequency component, obtained by downsampling and blurring the original image, and a residual/high-frequency component given by subtracting the upsampled low frequency component from the original image.
@@ 5.3. A Hierarchical Model
Planetary terrain spans several orders of magnitude in scale, from continental structure to meter-level detail, making one-pass generation infeasible. Several previous works (Du et al., 2024; Sharma et al., 2024; Zhang et al., 2025; Lee et al., 2023; Zhou and Tang, 2024) have shown that MultiDiffusion produces repetitive results when poorly conditioned, a problem that InfiniteDiffusion inherits. We therefore organize generation into a small hierarchy of models operating at progressively finer resolutions. Each stage refines and conditions on the one above, maintaining large-scale coherence while producing realistic local detail. All models share a common EDM2 (Karras et al., 2024b) backbone with the modifications proposed in sCM (Lu and Song, 2025). The hierarchy begins with a coarse planetary model, which generates the basic structure of the world from a rough, procedural or user-provided layout. The next stage is the core latent diffusion model (Rombach et al., 2021), which transforms that rough structure into realistic 46km tiles in latent space. Finally, a consistency decoder expands these latents into a high-fidelity elevation map. All models are trained on random crops to learn a nearly translation-invariant representation, reflecting the fact that generation should be independent of absolute location. We visualize the coarse-to-fine pipeline in Figure  3.
@@
The core latent diffusion model synthesizes $512\times 512$ patches at a 90m resolution in signed-sqrt space, corresponding to 46km tiles. It predicts a 64x64 low-frequency elevation channel and latent map that compactly represents the corresponding $512\times 512$ image. To supply the latent codes, we train a separate VAE-style autoencoder that shares the same U-Net backbone but omits diffusion-specific components and skip connections. The model is optimized using L1 and LPIPS losses with a weak KL term to prevent overfitting. After training, the encoder processes each 2048×2048 tile in the dataset as a whole, and the resulting latent image is precomputed and stored alongside the tile. To maximize local quality, the imprecise VAE decoder is discarded, and a diffusion decoder learns to expand these latents into realistic and high-resolution residuals. Conditioning in the diffusion decoder is implemented by concatenating nearest-neighbor interpolated latents to the noisy input image at each diffusion step.
@@
To facilitate long-range global coherence, the core model is conditioned on $4\times 4$ patches of elevation data. Each pixel of the patch is about 23km, with the model prediction corresponding to the $2\times 2$ interior. Each patch contains 3 channels: the mean elevation of the pixel, the 5th percentile elevation of the pixel, and a binary mask indicating which pixels have data available. We also provide the model with the tile’s mean temperature, temperature variation, annual precipitation, and precipitation seasonality for additional coherence and user control. Since climatic data is not available in the ocean, we replace missing values with a standard gaussian, ensuring the model accepts any combination of climatic variables in ocean regions. These values are injected into the model as a flattened vector alongside the noise embedding.
@@ 5.4. Real-Time Planetary Scale Synthesis
The coarse diffusion model facilitates this hierarchy by producing the conditioning variables required by the core latent model. The user provides initial maps for this model using hand-drawn sketches or procedural noise, which we found works as well as purely learned methods while providing additional control. During inference, these inputs are corrupted with gaussian noise according to the user’s preference on a per-channel basis, and concatenated against the usual diffusion inputs. The model follows the EDM2 design but with no downsampling or upsampling operations. This limits the receptive field by design, preventing the model from drifting toward the massive continental structures present in Earth data and avoiding conflicts where global priors override user guidance while still allowing strong local corrections.
@@ 6. Results
We evaluate Terrain Diffusion on visual fidelity and latency. Latency experiments use a single NVIDIA RTX 3090 Ti GPU. InfiniteDiffusion uses $T=2$ and a stride of 32 in the core model, and $T=1$ with a batch size of 1 in the coarse and decoder models. We use a separable linear weight window that is 1 at the window center and decreases linearly to a small value $\epsilon$ at the boundary, which reduces FID from $19.32$ with a constant map to $14.78$. The decoder model uses windows of size 512 with strides 384. Batch size is 16 for the core model and 1 for other components.
@@ 7. Discussion
The main limitation of our method arises from data scarcity at the coarsest levels of the hierarchy. Each additional level operates on a downsampled representation of the dataset, rapidly reducing the number of independent training examples available to learn large-scale structure. This can make it difficult to train a diffusion model with sufficient receptive field to enforce global coherence. This limitation is application-dependent. For natural terrain, meaningful global structure exists only at the scale of a small number of continental landmasses, making data scarcity unavoidable. In our case, externally specified procedural priors, such as Perlin noise, provide a practical source of coarse structure that is seed-consistent and controllable, while the learned hierarchy refines this signal at higher resolutions where data is abundant. In contrast, domains with effectively unbounded or tiling structure, such as textures, may support an entirely learned hierarchy.